Design a three-channel inverting adder to satisfy the following relation:
Vo = -(10V1-0.5dV2÷dt +2 V3)
1
Expert's answer
2019-03-01T11:29:11-0500

The scheme is presented below:



and its output signal


"V_{\text{out}}=-10V_1+0.5\frac{dV_2}{dt}-2V_3"

can be written this way:


"V_{\text{out}}=-\frac{R_\text{f}}{R_1}V_1+R_\text{f}C\frac{dV_2}{dt}-\frac{R_\text{f}}{R_3}V_3,"

so all you have to do is choose the resistors and capacitor satisfying these conditions:


"R_\text{f}=10R_1=2R_3=0.5/C"